A delivery truck is 5km away from a car. The two started at the same time and headed towards the same direction. The delivery truck accelerates at 3m/s^2. If the car should overtake the truck in ten minutes, what should be the acceleration of the car?


1
Expert's answer
2021-05-07T09:44:58-0400

Let's first write the displacements of the car and delivery truck:


"x_{car}=\dfrac{1}{2}a_{car}t^2,""x_{truck}=x_0+\dfrac{1}{2}a_{truck}t^2."


At the time when the car overtakes the truck, "x_{car}=x_{truck}":


"\dfrac{1}{2}a_{car}t^2=x_0+\dfrac{1}{2}a_{truck}t^2,""a_{car}=\dfrac{2(x_0+\dfrac{1}{2}a_{truck}t^2)}{t^2},""a_{car}=\dfrac{2\cdot(5000\ m+\dfrac{1}{2}\cdot3\ \dfrac{m}{s^2}\cdot(600\ s)^2)}{(600\ s)^2}=3.03\ \dfrac{m}{s^2}."